# NOWA: Null-space Optical Watermark for Invisible Capture Fingerprinting and Tamper Localization

### v2

@@ Abstract
Ensuring the authenticity and ownership of digital images is increasingly challenging as modern editing tools enable highly realistic forgeries. Existing image protection systems mainly rely on digital watermarking, which is susceptible to sophisticated digital attacks. To address this limitation, we propose a hybrid optical-digital framework that incorporates physical authentication cues during image formation and preserves them through a learned reconstruction process. At the optical level, a phase mask in the camera aperture produces a Null-space Optical Watermark (NOWA) that lies in the Null Space of the imaging operator and therefore remains invisible in the captured image. Then, a Null-Space Network (NSN) performs measurement-consistent reconstruction that delivers high-quality protected images while preserving the NOWA signature. The proposed design enables tamper localization by projecting the image onto the camera’s null space and detecting pixel-level inconsistencies. Our design preserves perceptual quality, resists common degradations such as compression, and establishes a structural security asymmetry: without access to the optical or NSN parameters, adversaries cannot forge the NOWA signature. Experiments with simulations and a prototype camera demonstrate competitive performance in terms of image quality preservation, and tamper localization accuracy compared to state-of-the-art digital watermarking and learning-based authentication methods.
@@ 1 Introduction
We live in an era dominated by digital media, where images and videos have become a primary medium for communication, documentation, and evidence. At the same time, the technological advances that facilitate this ubiquity have also enabled the creation, manipulation, and dissemination of digital images and videos with unprecedented ease, eroding trust in visual media. From generative editing tools to large-scale diffusion models, synthetic forgeries can now mimic reality with alarming fidelity, prompting the increasingly common question: Is this real or AI?. For instance, a deepfake video of Ukrainian President Volodymyr Zelenskyy in 2022, demonstrates how fake or doctored media artifacts can significantly degrade peoples’ trust in digital media today. Further, this highlight the urgent need for robust image and video forensic methods to detect and counteract media attacks.
@@
While robustness is desirable in copyright protection and media tracing, it can be counterproductive in security tasks that demand sensitivity to even minimal tampering. Fragile watermarking addresses this need by making the watermark highly sensitive to changes, supporting pixel-level integrity verification [67, 53, 44]. However, fragile watermarking remains largely underexplored in the optical domain. Recent efforts [4] using coded apertures begin to bridge this gap by embedding imperceptible optical signatures but exhibit a notably low image quality and only offer a binary authenticity verification without the ability to localize tampering.
@@
Here, we introduce a hybrid optical–digital framework that unifies the PM encoding with learned digital processing for end-to-end image protection. This integration allows our system to proactively embed optical authentication cues at the moment of capture while preserving the visual characteristics expected from end-consumer imaging devices. More precisely, our approach builds on the observation that the null space of the imaging operator defined by the PM contains information that is optically unmeasurable but mathematically defined. We exploit this property to embed a Null-space Optical Watermark (NOWA) that is invisible during capture and recoverable only through the camera’s forward model. To restore high quality protected images, we use a Null-Space Network (NSN) [52]. The NSN learns the inverse optical mapping with measurement consistency, ensuring reconstructions preserve the NOWA. This design creates a built-in security asymmetry: without the optical parameters and trained NSN parameters, the NOWA signature cannot be forged.
Furthermore, our hybrid design enables flexible and high-accuracy verification via a two-stage mechanism: isolating the unique optical signature through a null space projection of the designed camera and classifying tampering with a dedicated neural network. We evaluate our approach using different attack scenarios in simulations demonstrating competitive performance in terms of image quality preservation (PSNR, SSIM), and tamper localization accuracy compared to state-of-the-art digital watermarking and learning-based authentication methods. We translate our digital designed phase mask to hardware and demonstrate in real experiments high quality protected images.
@@ 2 Related Work
Optically Coded Capture. Several works explore embedding authentication cues directly in the optical domain. Ma et al. [36] introduce physical refractive objects (totems) whose captured regions act as coarse signatures for manipulation detection. Ptychography-based phase encoding enables optical watermarking using single-shot and non-mechanical setups [62, 37]. Naturally occurring chromatic aberrations can also serve as intrinsic encoders [40], revealing forged regions through aberration inconsistencies. Sensor-level fingerprints have been explored via pattern noise [35] and learned noise residuals [8], enabling device-specific forgery localization.
@@ 3 Method
We propose an authentication method that relies on the advantages of a physical and digital solution. Our proposed hybrid authentication system integrates optical front-end engineering with a Null-Space Network to form a differentiable end-to-end pipeline. The light propagation through the optical system is simulated using Fourier optics [16, 27, 61], enabling the joint optimization of both the optical parameters and the neural network for the authentication task. Mathematically, the proposed hybrid watermarked image is
@@
where $\mathbf{x}\in\mathbb{R}^{n}$ denotes the input scene, $g_{\phi}(\cdot)$ models the optical image formation process using a phase mask parameterized by $\phi$, $f_{\theta}(\cdot)$ is the reconstruction NSN with parameters $\theta$, responsible for processing the optical encoded images and finally obtain the protected image $\mathbf{x}_{p}\in\mathbb{R}^{n}$. Following reconstruction, the final stage of our pipeline is a verification mechanism that operates on the protected image $\mathbf{x}_{p}$ to detect and localize tampering at the pixel level. In the following, we present more details about the imaging model, NSN, and tamper detection and localization.
3.1 Imaging Model
Our optical system $g_{\phi}$ consists of a conventional imaging lens combined with a custom phase mask positioned at the pupil plane. The combined effect of the lens and the phase mask defines the system’s point spread function (PSF), which encodes a unique physical signature into every captured image. In the paraxial regime, the corresponding phase shift imparted by the convex lens thickness profile with focal length $f$ and wavelength-dependent refractive index $n(\lambda)$ produces the lens transmission function [17]:
@@ 3.1 Imaging Model
3.2 Null-Space Networks for Reconstruction and Authentication
In computational imaging, the measurement process in (9) can be modeled as
@@ 3.2 Null-Space Networks for Reconstruction and Authentication
$\mathcal{N}(\mathbf{A}_{\phi})=\{\,\mathbf{z}\mid\mathbf{A}_{\phi}\mathbf{z}=0\,\},$
is completely invisible to the sensor. We exploit this unobservable subspace to embed a Null-space Optical Watermark (NOWA) that is invisible in the optical encoded image and cannot be imitated without access to $\mathbf{A}_{\phi}$. However, the encoded images are not directly usable for final users, and any standard reconstruction algorithm would treat the NOWA as noise and remove it. Therefore, we use a Null-Space Network (NSN) [51] designed to both reconstruct an accurate image and recover lost information in $\mathcal{N}$. More precisely, given a regularized estimate $\hat{\mathbf{x}}_{r}=r(\mathbf{y})\in\mathcal{N}(\mathbf{A}_{\phi})^{\perp}$, obtained via a regularized inverse (e.g., pseudoinverse, Tikhonov, or Wiener deconvolution), the NSN recovers a protected image defined as
@@
where $\Pi_{\mathcal{N}}=\mathbf{I}-\mathbf{A}_{\phi}^{\dagger}\mathbf{A}_{\phi}$ is the orthogonal projector onto $\mathcal{N}$. This formulation enforces strict measurement consistency, $\mathbf{A}_{\phi}f_{\theta}(\hat{\mathbf{x}}_{r})=\mathbf{A}_{\phi}\hat{\mathbf{x}}_{r}=\mathbf{y}$, while allowing the learned network $\mathcal{U}_{\theta}$ to populate the unobservable components. In this sense, the null space provides a mathematically secure embedding channel: its energy is annihilated by $\mathbf{A}_{\phi}$ but recoverable through projection $\Pi_{\mathcal{N}}$. Any modification to $\mathbf{x}_{p}$ that is inconsistent with $\mathcal{N}(\mathbf{A}_{\phi})$ can be detected by re-projecting the image and analyzing the residual null-space statistics. This observation forms the foundation of our detection and localization pipeline, which is described in the following section.
3.3 Tamper Detection and Localization
The authentication stage of our framework operates entirely in the null-space domain defined by $\mathbf{A}_{\phi}$. Given a protected reconstruction $\mathbf{x}_{p}$, we extract a raw signature map by projecting it onto the null space of $\mathbf{A}_{\phi}$
@@ 4.4 Adversarial Robustness Evaluation
In this case, the adversary directly targets the distribution of protected images by training a generative model $\mathcal{G}_{2}$ that learns to map public unprotected data into the protected domain $\mathcal{G}_{2}:\mathcal{X}_{\text{u}}\rightarrow\mathcal{X}^{\prime}_{\text{p}}$. The generated images $\mathbf{x}^{\prime}_{p}=\mathcal{G}_{2}(\mathbf{x}_{\text{u}})$ are subsequently processed by the Image Verification network $d_{\psi}$. Attack 3: Blind Deconvolution. From camera outputs, the adversary also can attempt to estimate the system PSF via blind deconvolution techniques and apply it to arbitrary public images $\hat{\mathbf{y}}=\hat{\mathbf{A}}_{\phi}\mathbf{x}_{\text{u}}.$ The generated $\hat{\mathbf{y}}$ is then passed through the protection function $f_{\theta}$ (asumming is exposed) and the verification module $d_{\psi}$.
For $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, we employ a diffusion-based image-to-image translation model [45], and for blind deconvolution we use the method of Eboli et al. [10]. As shown in Fig. 4, all three attacks produce images that are visually similar to our system, however fail to reproduce the intrinsic null-space correlations imposed by the physical camera model. Specifically, for $1,000$ spoofed images per attack, our system achieves detection accuracies (F1 scores) of 0.901, 0.913, and 0.946 for attacks 1, 2, and 3, respectively, confirming the robustness of our hybrid optical–digital design against both learned and physics-based adversaries. Notably, even when the NSN is accessible to the attacker (Attacks 1 and 3), the absence of the authentic optical system prevents the generation of valid NOWAs. Likewise, the failure of Attack 2 shows that, without the optical encoding, the protected and natural image distributions remain statistically indistinguishable.
@@ 5 Real Captures Experiments
We verify the proposed framework in real-world capture experiment by translating the forward model $\mathbf{A}_{\phi}$ used in simulation to a physical optical prototype. The setup comprises a Canon EOS5DMarkIV DSLR coupled with a 50mm f/1.8 YONGNUO lens and the designed phase mask fabricated using two-photon lithography on a fused-silica substrate. The phase mask was placed in the aperture plane of the YONGNUO lens. Additional fabrication and optical setup details are provided in the Supplementary Material. For accurate protection under real imaging conditions, we fine-tune the Null-Space Network $f_{\theta}$ and the detector $d_{\psi}$ for 30 epochs with a learning rate of $1\times 10^{-5}$, using scenes from the FFHQ dataset rendered through the measured point spread functions (PSFs) of the real prototype. Captured images, fine-tuning details, and calibration including PSF measurement and alignment procedures, are provided in the Supplementary Material.

### v3

@@ Abstract
Ensuring the authenticity and ownership of digital images is increasingly challenging as modern editing tools enable highly realistic forgeries. Existing image protection systems mainly rely on digital watermarking, which is susceptible to sophisticated digital attacks. To address this limitation, we propose a hybrid optical-digital framework that incorporates physical authentication cues during image formation and preserves them through a learned reconstruction process. At the optical level, a phase mask in the camera aperture produces a Null-space Optical Watermark (NOWA) that lies in the Null Space of the imaging operator and therefore remains invisible in the captured image. Then, a Null-Space Network (NSN) performs measurement-consistent reconstruction that delivers high-quality protected images while preserving the NOWA signature. The proposed design enables tamper localization by projecting the image onto the camera’s null space and detecting pixel-level inconsistencies. Our design preserves perceptual quality, resists common degradations such as compression, and establishes a structural security asymmetry: without access to the optical or NSN parameters, adversaries cannot forge the NOWA signature. Experiments with simulations and a prototype camera demonstrate competitive performance in terms of image quality preservation and tamper localization accuracy compared to state-of-the-art digital watermarking and learning-based authentication methods.
@@ 1 Introduction
We live in an era dominated by digital media, where images and videos have become a primary medium for communication, documentation, and evidence. At the same time, the technological advances that facilitate this ubiquity have also enabled the creation, manipulation, and dissemination of digital images and videos with unprecedented ease, eroding trust in visual media. From generative editing tools to large-scale diffusion models, synthetic forgeries can now mimic reality with alarming fidelity, prompting the increasingly common question: Is this real or AI?. For instance, a deepfake video of Ukrainian President Volodymyr Zelenskyy in 2022 demonstrates how fake or doctored media artifacts can significantly degrade people’s trust in digital media today. Further, this highlights the urgent need for robust forensic methods for images and videos to detect and counteract media attacks.
@@
While robustness is desirable in copyright protection and media tracing, it can be counterproductive in security tasks that demand sensitivity to even minimal tampering. Fragile watermarking addresses this need by making the watermark highly sensitive to changes, supporting pixel-level integrity verification [67, 53, 44]. However, fragile watermarking remains largely underexplored in the optical domain. Recent efforts [4] using coded apertures begin to bridge this gap by embedding imperceptible optical signatures, but exhibit a notably low image quality and only offer a binary authenticity verification without the ability to localize tampering.
@@
Here, we introduce a hybrid optical–digital framework that unifies the PM encoding with learned digital processing for end-to-end image protection. This integration allows our system to proactively embed optical authentication cues at the moment of capture while preserving the visual characteristics expected from end-consumer imaging devices. More precisely, our approach builds on the observation that the null space of the imaging operator defined by the PM contains information that is optically unmeasurable but mathematically defined. We exploit this property to embed a Null-space Optical Watermark (NOWA) that is invisible during capture and recoverable only through the camera’s forward model. To restore high-quality protected images, we use a Null-Space Network (NSN) [52]. The NSN learns the inverse optical mapping with measurement consistency, ensuring reconstructions preserve the NOWA. This design creates a built-in security asymmetry: without the optical parameters and trained NSN parameters, the NOWA signature cannot be forged.
Furthermore, our hybrid design enables flexible and high-accuracy verification via a two-stage mechanism: isolating the unique optical signature through a null space projection of the designed camera and classifying tampering with a dedicated neural network. We evaluate our approach using different attack scenarios in simulations, demonstrating competitive tamper localization accuracy compared to state-of-the-art digital watermarking and learning-based authentication methods. We translate our digitally designed phase mask to hardware and demonstrate in real experiments high-quality protected images.
@@ 2 Related Work
Optically Coded Capture. Several works explore embedding authentication cues directly in the optical domain. Ma et al. [36] introduce physical refractive objects (totems) whose captured regions act as coarse signatures for manipulation detection. Ptychography-based phase encoding enables optical watermarking using single-shot and non-mechanical setups [62, 38]. Natural chromatic aberrations can also serve as intrinsic encoders [40], revealing forged regions through aberration inconsistencies. Sensor-level fingerprints have been explored via pattern noise [35] and learned noise residuals [8], enabling device-specific forgery localization.
@@ 3 Method
We propose an authentication method that leverages the advantages of a physical and digital solution. Our proposed hybrid authentication system integrates optical front-end engineering with a Null-Space Network to form a differentiable end-to-end pipeline. The light propagation through the optical system is simulated using Fourier optics [16, 27, 61], enabling the joint optimization of both the optical parameters and the neural network for the authentication task. Mathematically, the proposed hybrid watermarked image is
@@
where $\mathbf{x}\in\mathbb{R}^{n}$ denotes the input scene, $g_{\phi}(\cdot)$ models the optical image formation process using a phase mask parameterized by $\phi$, and $f_{\theta}(\cdot)$ is the reconstruction NSN with parameters $\theta$, responsible for processing the optical encoded images and finally obtain the protected image $\mathbf{x}_{p}\in\mathbb{R}^{n}$. Following reconstruction, the final stage of our pipeline is a verification mechanism that operates on the protected image $\mathbf{x}_{p}$ to detect and localize tampering at the pixel level. In the following, we present more details about the imaging model, NSN, and tamper detection and localization.
3.1 Preliminaries
Imaging Model.
Our optical system $g_{\phi}$ consists of a conventional imaging lens and a custom phase mask at the pupil plane. The combined effect of the lens and the phase mask defines the system’s point spread function (PSF), which encodes a unique physical signature into every captured image. In the paraxial regime, the corresponding phase shift imparted by the convex lens thickness profile with focal length $f$ and wavelength-dependent refractive index $n(\lambda)$ produces the lens transmission function [17]
@@ 3.1 Preliminaries
Null-Space Networks for Reconstruction.
The measurement process in (9) can be modeled as
@@
$\mathcal{N}(\mathbf{A}_{\phi})=\{\,\mathbf{z}\mid\mathbf{A}_{\phi}\mathbf{z}=0\,\},$ (12)
is completely invisible to the sensor. To address this, Null-Space Networks (NSNs) [51] have been designed to both reconstruct an accurate signal and the lost information in $\mathcal{N}(\mathbf{A}_{\phi})$. More precisely, given a regularized estimate $\hat{\mathbf{x}}_{r}=r(\mathbf{y})\in\mathcal{N}(\mathbf{A}_{\phi})^{\perp}$, obtained via a regularized inverse (e.g., pseudoinverse, Tikhonov, or Wiener deconvolution), the NSN recovers an image defined as
@@
where $\Pi_{\mathcal{N}}=\mathbf{I}-\mathbf{A}_{\phi}^{\dagger}\mathbf{A}_{\phi}$ is the orthogonal projector onto $\mathcal{N}$. This formulation enforces strict measurement consistency, $\mathbf{A}_{\phi}f_{\theta}(\hat{\mathbf{x}}_{r})=\mathbf{A}_{\phi}\hat{\mathbf{x}}_{r}=\mathbf{y}$, while allowing the learned network $\mathcal{U}_{\theta}$ to recover the unobservable components.
3.2 Null-space Optical Watermark (NOWA)
We propose to exploit the unobservable subspace $\mathcal{N}(\mathbf{A}_{\phi})$ as a secure channel to embed a Null-space Optical Watermark (NOWA) that cannot be imitated without access to $\mathbf{A}_{\phi}$. Unlike digital watermarks, the NOWA is physically encoded during image formation and remains invisible in the raw capture. However, the encoded images are not directly usable for final users, and any standard reconstruction algorithm would treat the NOWA as noise and remove it. To prevent this, we use the NSN framework to recover the protected image $\mathbf{x}_{p}$ while anchoring the NOWA within it. By leveraging measurement consistency, the null space provides a mathematically secure embedding: its energy is annihilated by $\mathbf{A}_{\phi}$ but recoverable through projection $\Pi_{\mathcal{N}}$. Any modification to $\mathbf{x}_{p}$ that is inconsistent with $\mathcal{N}(\mathbf{A}_{\phi})$ can be detected by re-projecting the image and analyzing the residual null-space statistics. This observation forms the foundation of our detection and localization pipeline.
Tamper Detection and Localization.
The authentication stage of our framework operates entirely in the null-space domain defined by $\mathbf{A}_{\phi}$. Given a protected reconstruction $\mathbf{x}_{p}$, we extract a raw signature map by projecting it onto $\mathcal{N}(\mathbf{A}_{\phi})$
@@ 4.4 Adversarial Robustness Evaluation
In this case, the adversary directly targets the distribution of protected images by training a generative model $\mathcal{G}_{2}$ that learns to map public unprotected data into the protected domain $\mathcal{G}_{2}:\mathcal{X}_{\text{u}}\rightarrow\mathcal{X}^{\prime}_{\text{p}}$. The generated images $\mathbf{x}^{\prime}_{p}=\mathcal{G}_{2}(\mathbf{x}_{\text{u}})$ are subsequently processed by the image verification network $d_{\psi}$. Attack 3: Blind Deconvolution. From camera outputs, the adversary can also attempt to estimate the system PSF via blind deconvolution techniques and apply it to arbitrary public images $\hat{\mathbf{y}}=\hat{\mathbf{A}}_{\phi}\mathbf{x}_{\text{u}}.$ The generated $\hat{\mathbf{y}}$ is then passed through the protection function $f_{\theta}$ (asumming is exposed) and the verification module $d_{\psi}$.
For $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, we employ a diffusion-based image-to-image translation model [45], and for blind deconvolution we use the method of Eboli et al. [10]. As shown in Fig. 4, all three attacks produce images that are visually similar to our system; however, they fail to reproduce the intrinsic null-space correlations imposed by the physical camera model. Specifically, for $1,000$ spoofed images per attack, our system achieves detection accuracies (F1 scores) of 0.901, 0.913, and 0.946 for attacks 1, 2, and 3, respectively, confirming the robustness of our hybrid optical–digital design against both learned and physics-based adversaries. Notably, even when the NSN is accessible to the attacker (Attacks 1 and 3), the absence of the authentic optical system prevents the generation of valid NOWAs. Likewise, the failure of Attack 2 shows that, without the optical encoding, the protected and natural image distributions remain statistically indistinguishable.
@@ 5 Real Captures Experiments
We verify the proposed framework in a real-world capture experiment by translating the forward model $\mathbf{A}_{\phi}$ used in simulation to a physical optical prototype. The setup comprises a Canon EOS5DMarkIV DSLR coupled with a 50mm f/1.8 YONGNUO lens and the designed phase mask fabricated using two-photon lithography on a fused-silica substrate. The phase mask was placed in the aperture plane of the YONGNUO lens. Additional fabrication and optical setup details are provided in the Supplementary Material. For accurate protection under real imaging conditions, we fine-tune the Null-Space Network $f_{\theta}$ and the detector $d_{\psi}$ for 30 epochs with a learning rate of $1\times 10^{-5}$, using scenes from the FFHQ dataset rendered through the measured point spread functions (PSFs) of the real prototype. Captured images, fine-tuning details, and calibration, including PSF measurement and alignment procedures, are provided in the Supplementary Material.
@@ 6 Discussion
Acknowledgments
This work was supported by NSF award IIS-2107313 and DoD ARO W911NF-24-2-0008. We thank Aniket Dashpute and Vivek Boominathan for helpful discussions. We thank Salman Khan for fabricating the phase mask.
@@ 10 Real-World Capture Procedure and Additional Analysis
Robustness against real-world camera imperfections
The phase mask induces a deterministic, structured PSF, whereas contaminants (e.g., dust) and minor aberrations introduce stochastic optical distortions. The null-space projection acts as a matched filter, isolating the watermark’s structured residuals from uncorrelated noise. However, in our real-world experiments, imperfect calibration or slight misalignment degrade null-space estimation that raises the energy baseline of the null space. Even after precise PSF calibration, the residual null-space energy $\|\Pi_{\mathcal{N}}(\hat{\mathbf{x}}_{r})\|$ increases from a baseline of $\sim 10^{-5}$ (not 0 due to numerical precision limits) in simulation to $\sim 10^{-4}$ in real-world captures. While this indicates minor spectral leakage of unwanted signals, this “real-world noise floor” remains significantly lower than the residuals induced by tampering. Consequently, the detector successfully differentiates between the stochastic artifacts of physical imperfections and the specific null-space violations of a forgery.